OpenStudy (anonymous):

2. Ball 2 with the mass of 30 kg is moving to the left at 10 m/s. With what velocity should Ball 1, with a mass of 10 kg, move to the right and collide with Ball 2 so that Ball 1 rebounds with a velocity of 30 m/s, and Ball 2 with a velocity of 10 m/s after the collision?

OpenStudy (anonymous):

okay so you know that momentum is conserved, and you can use the equation m1v1i +m2v2i = m1v1f +m2v2f, yeah?

OpenStudy (anonymous):

Yeah..

OpenStudy (anonymous):

I got 10V=300.

OpenStudy (anonymous):

you know that m2 = 30kg m1=10 kg v2i=-10m/s v1f=30 m/s v2f=10m/s so when you plug it in you get (10)*v1i + (30)(-10)=(10)(30)+(30)(10) and so 10v1i = 300 (like you said lol oops :) ) soo the velocity at which ball 1 should be initially travelling is 30 m/s.

OpenStudy (anonymous):

Oooh, alright! Thank you.

OpenStudy (anonymous):

no problem [: